# TongUI: Building Generalized GUI Agents by Learning from Multimodal Web Tutorials

### v2

@@ 4 Dataset Statistics
We provide four key statistics to show the diversity of the collected GUI-Net-1M dataset. We show the operating system distribution in  Fig. 4(a). Our dataset covers a diverse range of operating systems, including Windows, Android, iOS, MacOS, and Linux, ensuring a broad representation of GUI interactions across both desktop and mobile environments. The application distribution is shown in  Fig. 4(b). Our dataset covers a wide variety of application categories, ensuring a broad and diverse range of user tasks. We show the step distribution in  Fig. 4(c). Our dataset includes GUI interactions with varying step lengths, ranging from single-step actions to 9-step tasks. The distribution shows that shorter tasks (1-3 steps) are more frequent, while longer tasks gradually decrease in numbers. The higher proportion of trajectories of 1-3 steps is partly due to the trajectory splitting strategy described in Section 3.5. The action distribution is shown in  Fig. 4(d). Click and Tap are the most frequent, reflecting their central role in both desktop and mobile interactions. Other common actions include HotKey, Swipe, and Drag, covering various forms of user interactions.
Figure 5: Data flow of data filtering.
We also provide statistics on the data filtering flow in  Fig. 5. Throughout the data filtering pipeline, we observe a gradual reduction in data volume, reflecting the progressive refinement of the dataset. After Duplicate Tutorial Removal, 72.3% of the data remains. Following Content Filtering, this is further reduced to 58.5%. Finally, after applying Trajectory Filtering, 33.0% of the original data is retained. This step-wise filtering process helps to include high-quality and GUI-relevant interactions in the final dataset.
SeeClick-9.6B [4] 30.0 72.2 52.0 78.0 32.5 55.7 53.4
UGround-7B [7] 63.6 82.5 60.3 82.8 80.4 73.3 70.4
OmniParser-GPT-4V [23] 63.6 91.3 57.0 93.9 51.0 81.3 73.0
ShowUI-2B [18] 61.1 76.3 75.5 92.3 63.6 81.7 75.1
Qwen2.5-VL-3B $\dagger$ 30.7 66.0 40.2 70.0 51.5 69.1 56.5
Qwen2.5-VL-7B $\dagger$ 60.0 86.6 71.6 91.2 67.5 85.2 78.6
TongUI-3B 77.1 92.3 77.7 92.6 74.8 87.8 83.6
TongUI-7B 80.0 93.8 79.5 91.9 81.6 89.1 86.0
Model Mobile Desktop Web Avg
Text Icon/Widget Text Icon/Widget Text Icon/Widget
SeeClick [4] 78.4 50.7 70.1 29.3 55.2 32.5 55.1
OS-Atlas-4B [39] 87.2 59.7 72.7 46.4 85.9 63.1 71.9
OS-Atlas-7B [39] 95.2 75.8 90.7 63.6 90.6 77.3 84.1
Qwen2.5-VL-3B $\dagger$ 77.2 54.5 74.7 35.7 68.8 60.1 64.2
Qwen2.5-VL-7B $\dagger$ 95.5 81.5 90.2 65.0 90.2 70.4 84.0
TongUI-3B 94.4 79.6 92.8 75.0 87.6 77.8 85.5
TongUI-7B 93.1 81.5 96.4 82.9 90.2 84.7 88.7
Table 2: Results on ScreenSpot-V2.
Agent Model Development Creative CAD Scientific Office OS Avg
SeeClick [4] 0.6 0.0 0.3 1.0 0.0 0.6 2.5 0.0 1.9 3.5 0.0 2.0 1.1 0.0 0.9 2.8 0.0 1.5 1.8 0.0 1.1
OS-Atlas-4B [39] 7.1 0.0 3.7 3.0 1.4 2.3 2.0 0.0 1.5 9.0 5.5 7.5 5.1 3.8 4.8 5.6 0.0 3.1 5.0 1.7 3.7
ShowUI-2B [18] 16.9 1.4 9.4 9.1 0.0 5.3 2.5 0.0 1.9 13.2 7.3 10.6 15.3 7.5 13.5 10.3 2.2 6.6 10.8 2.6 7.7
CogAgent-18B [10] 14.9 0.7 8.0 9.6 0.0 5.6 7.1 3.1 6.1 22.2 1.8 13.4 13.0 0.0 10.0 5.6 0.0 3.1 12.0 0.8 7.7
Aria-UI [43] 16.2 0.0 8.4 23.7 2.1 14.7 7.6 1.6 6.1 27.1 6.4 18.1 20.3 1.9 16.1 4.7 0.0 2.6 17.1 2.0 11.3
UGround-7B [7] 26.6 2.1 14.7 27.3 2.8 17.0 14.2 1.6 11.1 31.9 2.7 19.3 31.6 11.3 27.0 17.8 0.0 9.7 25.0 2.8 16.5
OS-Atlas-7B [39] 33.1 1.4 17.7 28.8 2.8 17.9 12.2 4.7 10.3 37.5 7.3 24.4 33.9 5.7 27.4 27.1 4.5 16.8 28.1 4.0 18.9
Table 3: Results on ScreenSpot-Pro.
@@
PaLM2-CoT [48] 39.6
OmniParser [23] 48.3 57.8 51.6 77.4 52.9 57.7
SeeClick-9.6B [4] 54.0 73.7 57.6 66.4 54.9 59.3
ShowUI-2B [18] 63.9 77.5 66.6 72.5 69.7 70.0
@@
5 Experiments
5.1 Setting
Additional Training Data. In addition to GUI-Net-1M, we also refine some publicly available GUI datasets. For GUI grounding, we adopt ShowUI-website [18], ShowUI-desktop [18], and AMEX [2]. For GUI navigation, following the guidance of ShowUI, we use the training set from AITW [31], Mind2Web [5], MiniWob [11], and GUIAct [3]. We prompt GPT-4o-mini to generate thoughts for these datasets with textual guidance and visual spots on images, obtaining 256K data. We combine these data with GUI-Net-1M for training, totally 1M+256K $\approx$ 1.3M data. To make sure the thought is consistent with the action, we explicitly add vision prompts to images, such as a spot that indicates the location of the clicked button. In our text prompt, we instruct the model to generate a thought based on textual guidance of the task, description of the action, and vision prompts on the image.
Evaluation Data. We evaluate the TongUI agent on offline benchmarks: ScreenSpot [4], ScreenSpot-V2 [39], ScreenSpot-Pro [15], AITW [31], AndroidControl [16], Mind2Web [5], and Baidu Experience (we manually annotate 102 data points from the collected Baidu Experience data for testing). We also evaluate the TongUI agent on an online benchmark: MiniWob [34]. We use the same 35 tasks selected by SeeClick[4] to evaluate our model.
Training. We use Qwen2.5-VL-3B/7B as our base vision-language-action model. We train our model with a context window of 8192. We set the maximum number of previous observations as 2, which results in at most three images in a trajectory. We limit each image to 1350 vision tokens. Our training involves two stages. Following existing methods [18], we utilize the SFT data (training set) of the evaluation dataset to fine-tune the model. We apply LoRA on our model’s vision encoder and LLM parts with rank 16 and alpha 32. This setting leads to 0.5% of total parameters. We set the batch size per device to 4 and the gradient accumulation steps to 8. We set the learning rate as $41e-$ 4 and AdamW optimizer for training. The training process on 8$\times$×A100-80G takes roughly four days for a 3B model and seven days for a 7B model.
Compared Methods. We mainly compare the TongUI agent with open-source methods, including ShowUI [18], SeeClick [4], OS-Atlas [39], CogAgent [10], etc.
Additional experimental results, such as data scaling effect, are presented in the Supplementary Material.
5.2 Main Results
5.2.1 Grounding
In Tab. 1,  Tab. 2, and  Tab. 3, we show the zero-shot grounding performance of TongUI on ScreenSpot, ScreenSpot-V2, and ScreenSpot-pro, respectively. Similar to previous works [18], grounding on icon is much harder than grounding on text. The collected data leads to significant improvements on the baseline Qwen2.5-VL model. Compared with ShowUI, our method has about 5 % - 20 % improvements. These results demonstrate that GUI-Net-1M can indeed improve the grounding capability of GUI agents.
5.2.2 Offline Navigation
AITW. AITW is a dataset about navigation on Android. The comparisons are shown in Tab. 4. We observe that the fine-tuned TongUI-3B model achieves better performance compared to ShowUI-2B by 1.6% in overall success rate. This shows that the collected data also improves the reasoning and planning capabilities of GUI agents on the navigation tasks. The larger model, TongUI-7B, gains 1.7% improvement compared to TongUI-3B, which is consistent with common sense.
@@ 5.2.2 Offline Navigation
TongUI TongUI-3B 32.0 52.0
TongUI TongUI-7B 38.2 69.3
@@
AndroidControl. We also evaluate our performance in the AndroidControl benchmark in Tab. 5. As we do not include the AndroidControl training set, we compare our model with the zero-shot methods mentioned in previous work  [16]. TongUI-7B model achieves the best performance of low-level instructions(12.6% improvement).
Mind2Web. In Tab. 6, the results of web navigation show that TongUI achieves the best performance on all metrics. Notably, cross-task, cross-website, and cross-domain tasks are more difficult than conventional GUI tasks. This highlights that the collected data improves the generalization capability of GUI agents. We argue that the reason is that the collected data involves diverse applications and operating systems, improving the generalization capability. Compared to ShowUI-2B, TongUI-3B has obvious improvements, especially on Elem. Acc and Step SR. Compared with ShowUI-Qwen2.5-VL-3B that fine-tunes the Qwen2.5-VL-3B model using the ShowUI dataset, TongUI-3B also has better performance (3.2% improvements), showing the quality of our data.
@@
CogAgent [10] 22.4 53.0 17.6 18.4 42.4 13.4 20.6 42.0 15.5
MindAct [5] 55.1 75.7 52.0 42.0 65.2 38.9 42.1 66.5 39.6
OmniParser [23] 42.4 87.6 39.4 41.0 84.8 36.5 45.5 85.7 42.0
ShowUI-2B [18] 39.9 88.6 37.2 41.6 83.5 35.1 39.4 86.8 35.2
SeeClick-9.6B [4] 28.3 87.0 25.5 21.4 80.6 16.4 23.2 84.8 20.8
@@
Baidu Experience. The collected Baidu experience evaluation set includes tasks in Chinese, which evaluate the performance of GUI agents across different languages. Results are shown in Tab. 7. Compared with the baseline model, TongUI has improvements on both action accuracy and success rate, showing better generalization capability brought by the GUI-Net-1M dataset.
Model Action Acc. Success
ShowUI-2B [18] 65.7 14.7

### v3

@@ Introduction
We provide four key statistics to show the diversity of the collected GUI-Net-1M dataset. We show the operating system distribution in  fig. 2(a). Our dataset covers a diverse range of operating systems, including Windows, Android, iOS, MacOS, and Linux, ensuring a broad representation of GUI interactions across both desktop and mobile environments. We provide statistics on the data filtering flow in  fig. 2(b). Throughout the data filtering pipeline, we observe a gradual reduction in data volume, reflecting the progressive refinement of the dataset. After applying duplicate tutorial removal, content filtering, and trajectory filtering, 33.0% of the original data is retained. This three-step filtering process helps to include high-quality and GUI-relevant interactions in the final dataset. We show the step distribution in  fig. 2(c). Our dataset includes GUI interactions with varying step lengths, ranging from single-step actions to 9-step tasks. The distribution shows that shorter tasks (1-3 steps) are more frequent, while longer tasks gradually decrease in numbers. The higher proportion of trajectories of 1-3 steps is partly due to the trajectory splitting strategy described in Section 3.5. The action distribution is shown in  fig. 2(d). Click and Tap are the most frequent, reflecting their central role in both desktop and mobile interactions.
Agent Model ScreenSpot Avg ScreenSpot-V2 Avg ScreenSpot-Pro Avg
Text Icon Avg Text Icon Avg Text Icon Avg
GPT-4o 17.3 18.8 18.3 - - - 1.3 0.0 0.8
SeeClick-9B (cheng2024seeclick) 68.6 38.2 53.4 67.9 37.5 55.1 1.8 0.0 1.1
OS-Atlas-4B (wu2024atlas) 86.1 62.1 76.8 81.9 56.4 71.9 5.0 1.7 3.7
OS-Atlas-7B (wu2024atlas) 92.2 75.1 85.1 92.2 72.2 84.1 28.1 4.0 18.9
ShowUI-2B (lin2024showui) 83.4 66.7 75.1 - - - 10.8 2.6 7.7
CogAgent-18B (hong2024cogagent) - - - - - - 12.0 0.8 7.7
Aria-UI (yang2024aria) 90.7 71.4 82.4 - - - 17.1 2.0 11.3
AGUVIS-7B (xu2024aguvis) 92.6 73.3 84.4 - - - - - -
UGround-7B (gou2024navigating) 79.5 68.1 70.4 - - - 25.0 2.8 16.5
UI-TARS-2B (qin2025ui) 89.3 73.0 82.3 91.0 75.3 84.7 39.6 8.4 27.7
UI-TARS-7B (qin2025ui) 93.5 84.8 89.5 95.3 86.4 91.6 47.8 16.2 35.7
UI-TARS-72B (qin2025ui) 91.1 85.4 88.4 92.5 87.3 90.3 50.9 17.5 38.1
Qwen2.5-VL-3B $\dagger$ 68.4 40.8 56.5 73.6 50.1 64.2 9.1 3.3 6.9
Qwen2.5-VL-7B $\dagger$ 87.7 66.4 78.6 92.0 72.3 84.0 17.4 4.5 12.5
TongUI-3B 90.9 76.5 83.6 91.6 77.5 85.5 26.4 4.1 18.0
TongUI-7B 91.6 80.4 86.0 93.2 83.0 88.7 35.1 8.0 24.7
TongUI-32B 94.1 82.9 88.5 95.8 86.7 92.1 45.9 12.6 33.1
Table 2: Results on ScreenSpot, and ScreenSpot-V2, and ScreenSpot-Pro. $\dagger$ means the results are reproduced. The best method is marked in bold, and the second-best method is underlined.
@@
PaLM2-CoT (zhang2024you) 39.6
OmniParser (lu2024omniparser) 48.3 57.8 51.6 77.4 52.9 57.7
SeeClick-9.6B (cheng2024seeclick) 54.0 73.7 57.6 66.4 54.9 59.3
ShowUI-2B (lin2024showui) 63.9 77.5 66.6 72.5 69.7 70.0
@@
Experiments
Setting
Evaluation. We use Qwen2.5-VL-3B/7B/32B as our base VLM model. We evaluate the TongUI agent on offline benchmarks: ScreenSpot (cheng2024seeclick), ScreenSpot-V2 (wu2024atlas), ScreenSpot-Pro (li2025screenspot), AITW (rawles2023androidinthewild), AndroidControl (li2025effects), Mind2Web (deng2023mind2web), and UI-Vision (nayak2025ui). We also evaluate the TongUI agent on an online benchmark: MiniWob (shi2017world), following the same setting in SeeClick(cheng2024seeclick).
Compared Methods. We mainly compare the TongUI agent with open-source methods, including ShowUI (lin2024showui), SeeClick (cheng2024seeclick), OS-Atlas (wu2024atlas), CogAgent (hong2024cogagent), AGUVIS (xu2024aguvis), AgentTrek (xu2024agenttrek), etc. We also compare TongUI with UI-TARS (qin2025ui) that only releases models without training details and training data. In contrast, we have fully open-sourced our data, code, and models.
Main Results
Grounding
In table 2, we show the zero-shot grounding performance of TongUI on ScreenSpot, ScreenSpot-V2, and ScreenSpot-pro. Similar to previous works (lin2024showui), grounding on icons is much harder than grounding on text. The collected data leads to significant improvements on the baseline Qwen2.5-VL model. Compared with ShowUI, our method has about 5 % - 20 % improvements. TongUI has a comparable performance to UI-TARS, but UI-TARS only releases the models without training details and data. These results demonstrate that GUI-Net-1M can indeed improve the grounding capability of GUI agents.
@@
AGUVIS AGUVIS-7B 61.5 80.5
AGUVIS AGUVIS-72B 66.4 84.4
UI-TARS UI-TARS-2B 68.9 89.3
UI-TARS UI-TARS-7B 72.5 90.8
TongUI TongUI-3B 73.3 91.5
TongUI TongUI-7B 76.0 91.9
@@
Model Basic Functional Spatial
GPT-4o 1.6 1.5 1.0
Geimni-Flash-2.0 0.5 0.4 0.1
Claude-3.7-Sonnet 9.5 7.7 7.6
ShowUI-2B (lin2024showui) 8.1 7.7 2.1
AriaUI25-3B (yang2024aria) 12.2 14.0 4.0
UGround-7B (gounavigating) 15.4 17.1 6.3
AGUVIS-7B (xu2024aguvis) 17.8 18.3 5.1
CogAgent-9B (hong2024cogagent) 12.0 12.2 2.6
UI-TARS-7B (qin2025ui) 20.1 24.3 8.4
Qwen2.5-VL-7B 1.2 0.8 0.5
TongUI-3B 22.4 17.4 6.5
TongUI-7B 24.4 22.5 7.2
TongUI-32B 24.5 24.8 11.3
Table 5: Results on UI-Vision
@@
CogAgent (hong2024cogagent) 22.4 53.0 17.6 18.4 42.4 13.4 20.6 42.0 15.5
MindAct (deng2023mind2web) 55.1 75.7 52.0 42.0 65.2 38.9 42.1 66.5 39.6
OmniParser (lu2024omniparser) 42.4 87.6 39.4 41.0 84.8 36.5 45.5 85.7 42.0
ShowUI-2B (lin2024showui) 39.9 88.6 37.2 41.6 83.5 35.1 39.4 86.8 35.2
SeeClick-9.6B (cheng2024seeclick) 28.3 87.0 25.5 21.4 80.6 16.4 23.2 84.8 20.8
AgentTrek (xu2024agenttrek) 45.5 84.9 40.9 40.8 82.8 35.1 48.6 84.1 42.1
UI-TARS-2B (qin2025ui) 62.3 90.0 56.3 58.5 87.2 50.8 58.8 89.6 52.3
UI-TARS-7B (qin2025ui) 73.1 92.2 67.1 68.2 90.9 61.7 66.6 90.9 60.5
UI-TARS-72B (qin2025ui) 74.7 92.5 68.6 72.4 91.2 63.5 68.9 91.8 62.1
@@
Offline Navigation
We evaluate the offline navigation capability of GUI agents on the AITW, AndroidControl, Mind2Web, and UI-Vision datasets, and results are shown in table 3table 4table 6, and table 5, respectively. No matter whether on a small model (3B) or a big model (32B), using GUI-Net-1M data leads to competitive performance. For example, TongUI-3B achieves better performance compared to ShowUI-2B by 1.6% on AITW and more than 10% on UI-Vision. The larger model, TongUI-7B, gains better performance compared to TongUI-3B, which is consistent with common sense. On the AndroidControl and UI-Vision datasets, TongUI achieves even better performance than UI-TARS. This highlights that the collected data improves the generalization capability of GUI agents. We argue that the reason is that the collected data involves diverse applications and operating systems, improving generalization.
Online Navigation
We evaluate the online navigation performance on MiniWob, as shown in table 7. Compared with offline navigation, online navigation is more challenging, since it processes dynamic environments, handles unexpected obstacles, and adapts to changes in the navigation path. In this case, TongUI, which learns from multimodal tutorials, achieves improvements again. Considering that the multimodal tutorials are offline data, the improvements confirm the generalization of TongUI.
Model Finetuned Score
CC-Net(SL) (humphreys2022data) $\checkmark$ 23.4
Pix2Act (shaw2023pixels) $\checkmark$ 55.2
AGUVIS-72B (xu2024aguvis) $\checkmark$ 66.0
SeeClick-9.6B (cheng2024seeclick) $\checkmark$ 67.0
Qwen2-VL-2B (wang2024qwen2) $\checkmark$ 66.8
ShowUI-2B (lin2024showui) $\checkmark$ 71.5
Qwen2.5-VL-3B $\times$ 0.3
Qwen2.5-VL-3B-ShowUI $\checkmark$ 67.7
TongUI-3B $\checkmark$ 72.7
TongUI-7B $\checkmark$ 73.9
TongUI-32B $\checkmark$ 74.3
Table 7: Results on MiniWob
User Study for Data Quality
To validate the effectiveness of our data filtering strategy and evaluate the data quality of the collected GUI-Net-1M dataset, we conduct two user studies. In both studies, five participants with substantial experience in GUI agent research (but not involved in our project) are asked to rate the quality of each data point on a scale from 0 (very poor) to 5 (excellent). As shown in fig. 3(a), we evaluate the impact of our Trajectory Filtering step. Specifically, we randomly sample 100 data points before filtering and 100 after filtering, and mix them. As shown in fig. 3(a), the average score increases from 3.22 to 4.12 after filtering, demonstrating that Trajectory Filtering significantly improves data quality by effectively removing low-quality or GUI-irrelevant steps. We also randomly sample 100 data points from the ShowUI dataset for comparison. The results are shown in fig. 3(b), and the ShowUI dataset receives an average score of 4.26. This suggests that data quality in GUI-Net-1M is comparable to that in the ShowUI dataset, while our data leads to further improvements.
Figure 3: Average scores from humans
Conclusion
In this paper, we present the TongUI framework that converts multimodal web tutorials for an 1M trajectory dataset for GUI agents. By defining suitable and rich sources, we can crawl multimodal tutorials about diverse applications in different operating systems. The proposed tutorial processing method can extract tasks and trajectories from tutorials, and we obtain the GUI-Net-1M dataset. Using this dataset, the tuned Qwen2.5-VL model achieves improvements on multiple commonly used benchmarks, demonstrating the effectiveness of the TongUI framework and the collected GUI-Net-1M dataset. In this method, we have to collect all tutorials and train the model once, which may overlook the potential for continual learning. In the future, we will explore the continual learning capability of GUI agents, enabling the agents to better adapt to new environments.
Acknowledgments
This work was supported by National Science and Technology Major Project (2022ZD0114900), the Natural Science Foundation of China (NSFC) under Grants (No. 62406009), and Opening Project of the State Key Laboratory of General Artificial Intelligence (SKLAGI2024OP01, SKLAGI2024OP14).
Figure 4: TongUI crawls and processes multimodal web tutorials from diverse sources into GUI trajectories. The collected GUI-Net-1M dataset spans over 280 applications across five operating systems, improving generalization of GUI agents.
Appendix A Details
Evaluation Metric.
After training, we perform a zero-shot evaluation of the model’s grounding capability on ScreenSpot, ScreenSpot-V2, and ScreenSpot-Pro. The annotation of these grounding benchmarks gives a bounding box for each interactive element. The action from the model is considered correct when the point predicted falls within the bounding box. For online evaluations, the environment of MiniWob provides reward functions, which suggests if a trajectory is correct. For offline evaluations such as Mind2Web, AITW, Baidu Experience, AndroidControl, and GUI Odyssey, the model will be asked to output actions with parameters. We evaluate if the action is correct by the exact match. For parameters, we do the exact match for string parameters and check if a point parameter falls within the bounding box based on annotations. The step is correct if and only if both of the action prediction and parameter prediction are correct.
Model Architecture
The model architecture is shown in fig. 5.
Figure 5: Illustration for TongUI model architecture.
Action Space
The used action space is shown in table 8.
Environment Action Definition
Desktop Click(x,y) Left single click on an element.
Input(value, x, y) Type a string into certain element.
Scroll(direction, x, y) Scroll screen with left/right/up/down direction. X and y is the start point.
LeftClickDouble(x, y) Left double click on an element
RightClickSingle(x, y) Right single click on an element
Drag(x1,y1,x2, y2) Drag the cursor to the position with left key pressed.
HotKey(value) Press the hot key.
Hover(x, y) Hover over an element.
Finish() Finish the task.
Mobile Tap(x, y) Finger tap on an element.
Input(value, x, y) Type a string into an element.
Swipe(x1, y1, x2, y2) Swipe the screen from the start position to the end position.
LongPress(duration, x, y) Press on an element with certain seconds.
PressHome() Press the home button.
PressBack() Press the back button.
Finish() Finish the task.
Table 8: Action space in trajectory generation.
Resource Consumption
[…]
Prompt
In the following, we show the system prompt in our agent framework.
Listing 1: System prompt for GUI-Net Agent
You are an assistant trained to navigate the computer/mobile screen.
Given a task instruction, a screen observation, and an action history sequence, output the next action and wait for the next observation.
Here is the action space:
{action_space}
Format your response as
Thought: <your reasoning process>
Action: <the next action>
Format the action as a JSON object with the following keys:
{"action": "ACTION_TYPE", "value": "element", "position": [x,y]}
You can output multiple actions at once, and use JSON array to represent multiple actions.
If value or position is not applicable, set it as None‘.
Position might be [[x1,y1], [x2,y2]] if the action requires a start and end position.
Position represents the relative coordinates on the screenshot and should be scaled to a range of 0-1.
In the following, we show a typical action space prompt for desktop.
Listing 2: Action space for GUI-Net Agent
1. CLICK‘: Click on an element, value is not applicable and the position [x,y] is required.
2. INPUT‘: Type a string into an element, value is a string to type and the position [x,y] is required.
3. SCROLL‘: Scroll the screen, value is the direction to scroll and the position is start position of the scroll operation.
4. LEFT_CLICK_DOUBLE‘: Left click on an element twice, value is not applicable and the position [x,y] is required.
5. RIGHT_CLICK_SINGLE‘: Right click on an element once, value is not applicable and the position [x,y] is required.
6. DRAG‘: Drag the cursor to the specified position with the left button pressed. Value is not applicable and position [[x1,y1], [x2,y2]] is the start and end position of the drag operation.
7. HOT_KEY‘: Press a hot key, value is the hot key and the position is not applicable.
8. WAIT‘: Wait for 5 seconds, and take a screenshot to check for any changes. Value and position are not applicable.
9. FINISH‘: Finish the task. Value and position are not applicable.
Appendix B Dataset
Applications
In GUI-Net, the collected data spans about 280 applications that are shown in figs. 6 and 7. The distribution of application category is shown in  fig. 8. Our dataset covers a wide variety of application categories, ensuring a broad and diverse range of user tasks.
Figure 6: Collected applications.
Figure 7: Collected applications.
Figure 8: Categories of collected applications.
SeeClick-9.6B (cheng2024seeclick) 30.0 72.2 52.0 78.0 32.5 55.7 53.4
UGround-7B (gou2024navigating) 63.6 82.5 60.3 82.8 80.4 73.3 70.4
OmniParser-GPT-4V (lu2024omniparser) 63.6 91.3 57.0 93.9 51.0 81.3 73.0
ShowUI-2B (lin2024showui) 61.1 76.3 75.5 92.3 63.6 81.7 75.1
AgentTrek (xu2024agenttrek) - - - - 51.5 81.7 -
AGUVIS-7B (xu2024aguvis) 67.1 93.8 77.7 95.6 75.2 88.3 84.4
UI-TARS-2B (qin2025ui) 68.6 90.7 75.5 93.0 74.8 84.3 82.3
UI-TARS-7B (qin2025ui) 85.7 95.9 85.2 94.5 83.5 90.0 89.5
UI-TARS-72B (qin2025ui) 82.5 94.9 88.6 89.7 85.0 88.7 88.4
Qwen2.5-VL-3B $\dagger$ 30.7 66.0 40.2 70.0 51.5 69.1 56.5
Qwen2.5-VL-7B $\dagger$ 60.0 86.6 71.6 91.2 67.5 85.2 78.6
TongUI-3B 77.1 92.3 77.7 92.6 74.8 87.8 83.6
TongUI-7B 80.0 93.8 79.5 91.9 81.6 89.1 86.0
TongUI-32B 80.0 94.8 84.3 96.3 84.5 91.3 88.5
Model Mobile Desktop Web Avg
Text Icon/Widget Text Icon/Widget Text Icon/Widget
SeeClick (cheng2024seeclick) 78.4 50.7 70.1 29.3 55.2 32.5 55.1
OS-Atlas-4B (wu2024atlas) 87.2 59.7 72.7 46.4 85.9 63.1 71.9
OS-Atlas-7B (wu2024atlas) 95.2 75.8 90.7 63.6 90.6 77.3 84.1
UI-TARS-2B (qin2025ui) 95.2 79.1 90.7 68.6 87.2 78.3 84.7
UI-TARS-7B (qin2025ui) 96.9 89.1 95.4 85.0 93.6 85.2 91.6
UI-TARS-72B (qin2025ui) 94.8 86.3 91.2 87.9 91.5 87.7 90.3
Qwen2.5-VL-3B $\dagger$ 77.2 54.5 74.7 35.7 68.8 60.1 64.2
Qwen2.5-VL-7B $\dagger$ 95.5 81.5 90.2 65.0 90.2 70.4 84.0
TongUI-3B 94.4 79.6 92.8 75.0 87.6 77.8 85.5
TongUI-7B 93.1 81.5 96.4 82.9 90.2 84.7 88.7
TongUI-32B 97.9 86.7 96.9 86.7 92.7 86.7 92.1
Table 10: Results on ScreenSpot-V2.
Agent Model Development Creative CAD Scientific Office OS Avg
SeeClick (cheng2024seeclick) 0.6 0.0 0.3 1.0 0.0 0.6 2.5 0.0 1.9 3.5 0.0 2.0 1.1 0.0 0.9 2.8 0.0 1.5 1.8 0.0 1.1
OS-Atlas-4B (wu2024atlas) 7.1 0.0 3.7 3.0 1.4 2.3 2.0 0.0 1.5 9.0 5.5 7.5 5.1 3.8 4.8 5.6 0.0 3.1 5.0 1.7 3.7
ShowUI-2B (lin2024showui) 16.9 1.4 9.4 9.1 0.0 5.3 2.5 0.0 1.9 13.2 7.3 10.6 15.3 7.5 13.5 10.3 2.2 6.6 10.8 2.6 7.7
CogAgent-18B (hong2024cogagent) 14.9 0.7 8.0 9.6 0.0 5.6 7.1 3.1 6.1 22.2 1.8 13.4 13.0 0.0 10.0 5.6 0.0 3.1 12.0 0.8 7.7
Aria-UI (yang2024aria) 16.2 0.0 8.4 23.7 2.1 14.7 7.6 1.6 6.1 27.1 6.4 18.1 20.3 1.9 16.1 4.7 0.0 2.6 17.1 2.0 11.3
UGround-7B (gou2024navigating) 26.6 2.1 14.7 27.3 2.8 17.0 14.2 1.6 11.1 31.9 2.7 19.3 31.6 11.3 27.0 17.8 0.0 9.7 25.0 2.8 16.5
OS-Atlas-7B (wu2024atlas) 33.1 1.4 17.7 28.8 2.8 17.9 12.2 4.7 10.3 37.5 7.3 24.4 33.9 5.7 27.4 27.1 4.5 16.8 28.1 4.0 18.9
UI-TARS-2B (qin2025ui) $47.4$ $4.1$ $26.4$ $42.9$ $6.3$ $27.6$ $17.8$ $4.7$ $14.6$ $56.9$ $17.3$ $39.8$ $50.3$ $17.0$ $42.6$ $21.5$ $5.6$ $14.3$ $39.6$ $8.4$ $27.7$
UI-TARS-7B (qin2025ui) $58.4$ $12.4$ $36.1$ $50.0$ $9.1$ $32.8$ $\mathbf{20.8}$ $9.4$ $\mathbf{18.0}$ $63.9$ $\mathbf{31.8}$ $\mathbf{50.0}$ $\mathbf{63.3}$ $20.8$ $53.5$ $30.8$ $\mathbf{16.9}$ $24.5$ $47.8$ $16.2$ $35.7$
TongUI-32B 44.8 6.2 26.1 45.5 7.7 29.6 26.4 6.3 21.5 61.8 27.3 46.9 61.8 27.3 48.3 45.8 11.2 30.1 45.9 12.6 33.1
Table 11: Results on ScreenSpot-Pro.
Appendix C More Experiments
Data ScreenSpot Mind2Web
Task Website Domain
No SFT 56.5 0.4 1.0 1.7
Refined data 68.0 39.7 35.5 40.7
+ WikiHow 50K 75.8 42.1 39.6 44.4
+ Baidu 50K 78.7 43.4 41.6 45.5
+ Video 50K 79.6 44.2 42.6 46.0
+All data 83.6 48.8 48.1 49.5
Table 12: Performance scaling over data sources, where Step SR is reported for Mind2Web.
Grounding Details
We provide detailed results on ScreenSpot, ScreenSpot-V2, and ScreenSpot-Pro, as shown in tables 9, 10 and 11, respectively.
Setting Metric Run 1 Run 2 Run 3 Mean ± Std
Cross-Task Elem. Acc 52.0 52.9 52.5 52.5 ± 0.451
OP. F1 88.6 88.6 88.8 88.7 ± 0.115
Step SR 47.6 48.5 48.0 48.0 ± 0.451
Cross-Web Elem. Acc 52.7 51.8 52.4 52.3 ± 0.458
OP. F1 86.0 86.3 86.6 86.3 ± 0.300
Step SR 46.1 45.7 46.5 46.1 ± 0.400
Cross-Domain Elem. Acc 54.1 53.8 53.6 53.8 ± 0.252
OP. F1 88.7 88.5 88.4 88.5 ± 0.153
Step SR 49.7 49.6 49.4 49.6 ± 0.153
Table 13: Error bar on Mind2Web.
Method Model High Low
[…]
TongUI TongUI-3B 73.3 91.5
TongUI TongUI-7B 76.0 91.9
Table 14: Step accuracy on AndroidControl. Experiments marked with $\dagger$ is from  (li2025effects). “LT-10k-r64” means the model is tuned by LoRA with 10k training trajectories and rank 64.
Model FT? Step Accuracy
CogAgent (hong2024cogagent) No 11.8
[…]
TongUI-3B Yes 71.04
TongUI-7B Yes 74.81
Table 15: Step accuracy on GUI Odyssey. “FT?” indicates if the model is fine-tuned on GUI Odyssey training set. Experiments mared with $\dagger$ is reported by previous work (lu2024gui)
Model Action Accuracy Success Rate
ShowUI-2B(lin2024showui) 65.7 14.7